### v1

# Enantiomer-Selective Magnetoresistance in Chiral Gold Nanocrystals by Magnetic Control of Surface Potentials

Abstract
Chiral nanomaterials offer intriguing possibilities for novel electronic and chemical applications. Here, we report the discovery of an enantiomer-selective magnetoresistance effect in chiral gold nanocrystals. Based on precise control of nanocrystal chiral morphology using amino acid-directed synthesis, we demonstrate that an external magnetic field can dramatically modulate resistance in an enantiomer-specific manner. For a given enantiomer, a magnetic field in one direction alters the resistance by over an order of magnitude, while the opposite field direction leaves it unchanged. This asymmetric response reverses for the opposite enantiomer. We attribute this phenomenon to a novel chirality-driven charge trapping mechanism, where the interplay between the chiral nanocrystal morphology and the magnetic field selectively modifies the surface potential. The magnitude and sign of the magnetoresistance can be further tuned by the surface chemistry of the nanocrystal, as demonstrated through sulfide treatment. Our findings reveal a new form of chirality-dependent magnetoresistance, distinct from previously known effects such as chirality-induced spin selectivity and electric magnetochiral anisotropy. The ability to remotely control surface potentials of chiral nanostructures using magnetic fields could enable novel approaches in catalysis, drug delivery, and nanoelectronics.
1 Introduction
The exploration of chirality has unlocked exciting possibilities for developing unconventional nanomaterials with distinctive properties and diverse applications, including chiroplasmonics [1], chiral catalysis [2], spintronics [3], enantiomer discrimination [4], and enantiomer-dependent chemical and biological responses [5]. Chiral quantum materials are gaining rapidly increasing research interest due to their fascinating magneto-transport properties, such as the chirality-induced spin selectivity (CISS) [6, 7] and electric magnetochiral anisotropy (EMCA) [8].
In CISS, chiral organic molecules (e.g., DNA or peptides) are sandwiched between a ferromagnetic electrode and a metal, exhibiting dramatically different resistance upon flipping the electrode’s magnetization [9, 10, 11]. Although the mechanism of CISS magnetoresistance (MR) is still under debate [12, 13, 14, 15], the ferromagnetic electrode, rather than a magnetic field, is believed to be essential for generating observable MR. In EMCA, chiral inorganic or organic crystalline conductors [16, 17, 18, 19, 20, 21] exhibit current direction-dependent MR in an external magnetic field. Furthermore, EMCA adheres to Onsager’s reciprocal relation by giving the same zero-bias resistance at opposite magnetic fields, while CISS MR violates reciprocity by showing significantly different zero-bias resistance for opposite magnetizations in the electrode [22, 15].
The chirality commonly originates in the molecule or crystal structure in these studies. Recently, inorganic materials with chiral morphology, for example, chiral gold nanocrystals (Au NCs) [23, 24], have been synthesized in a large scale and garnered considerable attention because they bring novel modes for molecular recognition and biological regulation [25, 26]. For example, chiral Au NCs with strong chiroplasmonic optical activity can generate superchiral electromagnetic fields for enantioselective sensing of adsorbed amino acids and proteins [27]. The high dissymmetry factor of chiral Au NCs enables the design of efficient circularly polarized light-detecting transistors [28] and sensitive chiroplasmonic hydrogen sensors [29]. Furthermore, the chiral selectivity between chiral molecules and chiral Au NCs also results in distinguishable electrochemical behaviors for enantioselective catalysis and chiral discrimination [30, 31]. Notably, chiral Au NCs can regulate the maturation of immune cells and exhibit distinct immune responses by their chirality [5, 32]. Known the versatile surface chemistry, the magneto-transport of chiral Au NCs will provide deep insights into controlling these chirality-dependent processes by a magnetic field.
In this work, we discovered a unique chirality-driven MR in chiral Au NCs by applying an external magnetic field. We synthesized Au NCs with pinwheel-like chiral morphology by an amino-acid-directed strategy and measured their resistance using conducting atomic force microscopy (cAFM). Distinct from the known CISS or EMCA, the resistance of chiral Au NCs is largely enhanced (positive MR) in one field orientation and nearly unchanged in the opposite field, showing a salient asymmetric feature with respect to the zero-field case. The asymmetric order was reversed when we switched to NCs with opposite handedness, verifying the role of chirality. By modifying the NC surface capping layer with sulfide, we can also reduce the resistance (negative MR) by one field direction and keep it unchanged in the opposite field. We rationalize the usual MR by the magnetochiral-driven charge trapping effect [13] in the surface capping layer, which controls the surface potential and the total resistance. Our work provides valuable insights to understand chirality-driven transport effects and offers implications for selectively controlling the surface potential and manipulate the surface activity for chemical and biological reactions.
2 Synthesis and characterization of chiral Au NCs
Figure 1: Synthesis, characterization, and chiroptical properties of chiral Au NCs. (a) Schematic illustration of the processes of Cys-directed growth of chiral Au NCs with opposite chirality. The pyramidal subunits on the trisoctahedron intermediates are rotated clockwise or counterclockwise. (b) The g-factor spectra. The insets show corresponding geometrical models of the chiral Au NCs with CTAB bilayers. (c-n) SEM images and corresponding geometric models of individual chiral Au NCs synthesized with L-Cys (c, d, e, f, g, h) and D-Cys (i, j, k, l, m, n) as chiral inducers, respectively. SEM images and geometric models in (c, f, i, l), (d, g, j, m), and (e, h, k, n) are viewed from the $\braket{111}$, $\braket{100}$, and $\braket{110}$ axes of the corresponding NCs, respectively. Note that the CTAB bilayers were omitted from the models in (f-k) for clarity. (o, p) Low-magnification SEM images of chiral Au NCs synthesized from L-Cys (o) and D-Cys (p), respectively. The L-Au and D-Au NCs with mean sizes of $150.2\pm 4.8$ nm and $154.4\pm 3.6$ nm were synthesized with 80 $\mu L$ of seeds. Scale bars: 100 nm in (n) and (p). (c-e, l-n) share the same scale bar in (n). (o) shares the same scale bar in (p).
The Au NCs with intrinsic chirality were synthesized based on an amino-acid-directed chirality transfer strategy [23, 30] (see Methods and Figure 1). With L- or D-cysteine (Cys) as chiral inducers, spherical Au seeds were grown into chiral Au NCs with opposite chirality. After the syntheses, the Au NCs were collected by centrifugation and redispersed in water with cetyltrimethylammonium bromide (CTAB) as the stabilizing agent. Note that CTAB (achiral) is adsorbed on Au surfaces to form a core-shell structure [33]. The chiral Au NCs synthesized with L- and D-Cys exhibit inverted circular dichroism (CD) spectra with respect to each other, with dissymmetry factors (g-factors) of approximately -0.033 and 0.041, respectively (Figure 1b and Extended Data Fig. 1). From the scanning electron microscopy (SEM) images, the pyramid-like subunits on the chiral Au NCs are rotated clockwise or counterclockwise with respect to the achiral trisoctahedral Au NCs obtained with racemic Cys (Figures  1c,1f,1i,1l, and Extended Data Fig. 2). The rotation direction of the pyramidal subunits is dictated by the chirality of the chiral inducers, due to the enantioselective interactions of L- or D-Cys with R- or S- chiral Au surfaces [23, 30]. After the rotation, the four adjacent pyramids on the trisoctahedron no longer share one vertex, but instead form a pinwheel-like structure. As shown in Figures  1d and 1m, the pinwheel-like structure of the chiral Au NCs is rotated clockwise and counterclockwise, respectively. To clarify, the Au NCs obtained with L- and D-Cys are designated as L-Au and D-Au NCs, respectively, and the achiral Au NCs with racemic Cys is designated as A-Au NCs (Extended Data Table 1). Furthermore, by adjusting the amounts of seeds in the growth processes, chiral Au NCs can be synthesized with sizes ranging from 150.2 to 475.5 nm with tunable chiroplasmonic properties (Extended Data Figs. 3,  4, and Extended Data Table 2).
3 Magnetoresistance of chiral Au NCs
Figure 2: Magneto-transport and asymmetric MR for chiral Au NCs. (a) An AFM image of D-Au NCs on the indium tin oxide (ITO) substrate. The scale bar is 1 $\mu$m. (b) Schematic of the ITO/chiral Au NC/Pt junction constructed via cAFM. Take the D-Au NC as an example. A permanent magnet (300 Oe) is placed beneath the ITO substrate. (c) and (d) The experimental (dots) and fitted (line, according to Eq. (1)) current-voltage curves for the L-Au NC and D-Au NC under different directions of magnetic fields, where the south pole of the magnet is pointing perpendicularly toward (+B, red) and away from (–B, blue) the ITO plane. 0B (black) stands for zero magnetic field. (e) and (f) The MR of the L-Au NC and D-Au NC under different directions of an external magnetic field.
We measured the MR of the chiral Au NCs by cAFM with a Pt tip at room temperature. Using chiral Au NCs with size of $\sim$150 nm as exemplars, we dispersed L- and D-Au NCs on an indium tin oxide (ITO) substrate, enabling MR measurements at the single-nanoparticle level (Figure 2). The size of the L- and D-Au NC is $150.2\pm 4.8$ and $154.4\pm 3.6$ nm, respectively. We recorded current-voltage (I-V) relationships of individual chiral Au NCs via the ITO/chiral Au NC/Pt junction. An external magnetic field (B = 300 Oe) was applied orthogonally to the ITO substrate using a permanent magnet placed below the ITO (Figure 2b). Unlike CISS MR studies [34], our setup did not incorporate a magnetic substrate or tip. The insulating CTAB molecular shell on Au NCs contributes significantly more to the resistance than the metallic Au NC core [35], resulting in small currents of tens of nA at 1.0 V bias. Figures 2c and 2d present the averaged I-V curves for L- and D-Au NCs, respectively, derived from hundreds of individual measurements (see Extended Data Fig. 5). The I-V curves were measured 320–679 times on more than 30 individual Au NCs from at least three parallel samples. The error bar was calculated as the standard deviation of all results. We observed an asymmetric MR behavior between +B and –B (Figure 2). For L-Au NCs, the I-V curve at +B closely resembles the zero-field case within error bars, while at –B, the resistance increases dramatically by 20–50 times. Conversely, for D-Au NCs, –B barely affects the resistance, while +B increases it by $\sim$20 times. This reversal of asymmetric behavior between L- and D-Au NCs strongly indicates its chirality-dependent origin. We quantify MR as the relative change in resistance due to the B-field: MR=$[R(\textbf{B})-R(0)]/R(0)$. The results show that MR is negligible at +B (–B) but substantial at –B (+B) for L(D)-Au NCs, respectively (Figures 2e and 2f).
To determine whether the chiral morphology of the Au core or the chiral ligand shell is primarily responsible for the observed MR results, we conducted control experiments using three additional types of Au NCs. In addition to the L-Au and D-Au NCs, we investigated achiral Au NCs synthesized with racemic Cys (A-Au, Extended Data Fig.2), L-cysteine-modified achiral Au NCs (Lm-A-Au), and D-cysteine-modified achiral Au NCs (Dm-A-Au). We prepared Lm-A-Au and Dm-A-Au by modifying achiral non-Cys Au NCs with L- and D-Cys, respectively (Extended Data Figs.6-7). Our initial measurements of the I-V behavior of A-Au NCs revealed negligible MR for ±B (Figures 3a, 3d, and Extended Data Fig. 8). For Lm-A-Au and Dm-A-Au NCs, while the asymmetric order between ±B showed a trend similar to L- and D-Au NCs, the magnitude of their MR was significantly smaller (Figure 3). The I-V curves under ±B and zero-field conditions were only marginally or hardly distinguishable when considering error bars. These findings strongly suggest that the substantial MR observed in chiral Au NCs is primarily attributable to the chiral morphology of the NC core rather than the molecular chirality of the ligand shell. This conclusion highlights the crucial role of the NC’s intrinsic morphology chirality in determining its MR properties.
Figure 3: Magnetotransport of A-Au, Lm-A-Au, and Dm-A-Au NCs. (a)-(c) Current-voltage curves of the A-Au, Lm-A-Au, and Dm-A-Au NCs (from left to right) under magnetic field directions of +B (red), –B (blue), and 0B (black), respectively. At least 430 current-voltage curves were obtained for all samples, with approximately 10 measurements performed on each Au NC. The error bar was defined as the standard deviation of all current values measured at each magnetic field condition.
Having established the chiral origin of the observed MR, we proceeded to investigate the robustness of this chirality-driven effect under various conditions, including different NC sizes, magnetic field strengths, and AFM forces, to rule out potential artifacts. Using L-Au NCs as a model system, we observed that the current through chiral Au NCs increases gradually as the NC size increases from 150.2 to 350.5 nm, consistent with previously reported transport behaviors of metal nanoparticles [35] (Extended Data Figs. 9a and 10). Concurrently, the overall MR value decreases with increasing NC size, while the MR values under –B consistently remain significantly larger than those under +B (Extended Data Fig. 9d), confirming the persistence of asymmetric MR across different NC sizes. The magnitude of the asymmetric MR also varies proportionally with the applied magnetic field strength. For L-Au NCs, as the external magnetic field decreases from 300 to 3 Oe, the current remains nearly constant for +B, while it increases from 0.5 to 21 nA for –B at a bias of -1.0 V. This corresponds to a substantial MR drop from 5100% to 18% (Extended Data Figs. 9b, 9e, and 11). In cAFM measurements, the contact between the AFM tip and Au NCs significantly influences the total resistance [36, 37]. To exclude potential artifacts arising from this contact, we evaluated the MR of L-Au NCs under various AFM forces. As anticipated, the current values increase with incremental force due to the compression of the molecular layer and the increased contact area between the tip and the Au NC. Notably, the asymmetric MR remains robust across different applied forces, although the MR at –B exhibits a weakly nonmonotonic change with increasing force (Extended Data Figs. 9c and 9f). This robustness against varying AFM force may be attributed to the stable morphology of crystalline nanoparticles compared to organic materials. These comprehensive investigations across different experimental parameters underscore the intrinsic nature and reliability of the observed chirality-driven magnetoresistance effect in Au NCs.
4 Control the sign of magnetoresistance
Figure 4: Magnetoresistance characterization of the sulfide-modified chiral Au NCs. (a) and (b) The experimental (dots) and fitted (line) current-voltage curves of Sm-L-Au and Sm-D-Au NCs at the magnetic field directions of +B (red), –B (blue), and 0B (black). We measured 180–650 current-voltage curves from more than 20 individual Au NCs and estimated the error bar by the standard deviation. (c) and (d) The MR of Sm-L-Au and Sm-D-Au NCs changing with bias under different external magnetic field directions. The Sm-L-Au and Sm-D-Au NCs were prepared by treating the L-Au and D-Au NCs with sodium sulfide, respectively.
We discovered that the sign of asymmetric MR is sensitive to the chemical composition of the surface molecular layer on chiral Au NCs. We further treated the chiral Au NCs with sodium sulfide, resulting in additional sulfide adsorption into the CTAB layers on the Au NCs [38]. For clarity, we designate the sulfide-modified L- and D-Au NCs as Sm-L-Au and Sm-D-Au NCs, respectively. X-ray photoelectron spectroscopy studies confirm the presence of sulfide on their surfaces (Extended Data Fig. 12). SEM and spectra studies demonstrate that Sm-L-Au and Sm-D-Au NCs maintain similar chiral morphologies and optically chiral properties to their non-modified counterparts (Extended Data Figs. 13 and 14). Dynamic light scattering reveals slight size decreases for Sm-L-Au and Sm-D-Au NCs, attributed to the partial desorption of CTAB upon sulfide modification (Extended Data Fig. 15 and Extended Data Table 3).
We investigated the MR of Sm-L-Au and Sm-D-Au NCs using the same cAFM setup (Figure 4 and Extended Data Fig. 16). At zero magnetic field, sulfide-modified NCs exhibit significantly higher conductance than pristine NCs, likely due to the thinner CTAB molecular layer. Despite this conductance change, the asymmetric nature of MR persists. For Sm-L(D)-Au NCs, –B (+B) field dramatically reduces the resistance by several times, while +B (–B) barely affects it. In contrast to L(D)-Au NCs, Sm-L(D) NCs display negative MR, as shown in Figure 4. This sign change in MR suggests that the magnetotransport properties are intimately related to the characteristics of the molecular shell layer.
5 Discussions
5.1 Mechanisms on the potential modification.
The MR observed in chiral NCs exhibits unique characteristics that distinguish it from known magneto-transport phenomena. Firstly, it is independent of current direction and violates the general Onsager’s reciprocal relation. Secondly, its activation is triggered by a selected magnetic field direction, showing the asymmetric feature. These behaviors contrast sharply with established magneto-transport mechanisms: (i) In nonmagnetic metals, the ordinary MR, induced by the Lorentz force, is proportional to $\mathbf{B}^{2}$ and symmetric for ±$\pm\mathbf{B}$ [39]. (ii) In chiral conductors, EMCA produces an additional MR proportional to $\mathbf{B\cdot I}$, dependent on both current direction and amplitude [8]. (iii) The CISS effect typically yields MR that is symmetric between opposite electrode magnetizations (e.g. , Refs. [9, 40, 41, 42]) relative to the paramagnetic case. Our observed MR deviates from these established phenomena. Unlike EMCA, but similar to CISS, it violates Onsager’s relation, showing varied resistance at ±B in the zero-bias limit. However, it also differs from typical CISS MR in its asymmetry. Existing theoretical models proposed to explain CISS MR [43, 44, 45, 46, 47] fail to interpret our results. For instance, simple spin filter/polarizer models [43, 44, 48] only reproduce EMCA-like effects. More sophisticated models involving chirality-induced spin-transfer torques [45, 47] or vibration-enhanced spin-orbit coupling [46, 49] require magnetic electrodes and predict symmetric MR for opposite magnetizations. These models cannot account for the pronounced MR asymmetry we observe using non-magnetic leads. Given these discrepancies, we must conclude that our observed effect cannot be explained by ordinary MR, EMCA, or previously proposed CISS mechanisms. This necessitates the development of a new theoretical framework to understand the unique magneto-transport behavior in chiral NCs.
Figure 5: The magnetic field modified tunneling barrier for chiral Au NCs. (a)-(b) Schematics of the 1D tunneling barrier under 0B (gray region), +B (red dashed curve) and –B (blue dashed curve) for the L- and Sm-L-Au NCs, respectively. (c)-(d) Alignment of HOMO and LUMO of Au NCs. The energy level data was measured by UPS and LEIPS. The work functions of ITO and Pt are measured with respect to the vacuum level ($E_{\mathrm{VAC}}=0\,\text{eV}$).
Table 1: Fitted barrier heights ($U$) according to Eq. (1) at different magnetic fields (B) for L/D and Sm-L/D-Au NCs from the experimental data in Figures 2 and 4.
L D Sm-L Sm-D
0B 0.57 0.61 0.98 0.98
+B 0.58 1.15 0.97 0.56
-B 1.26 0.58 0.54 0.93
To understand the unconventional MR, we first adopted a barrier-modulation model to analyze the observed asymmetric MR. As mentioned above, the insulating molecular shell covered on the Au NC determines the main resistance. The molecular CTAB layer is $\sim$3 nm thick [33] and appears between the chiral Au NC and both electrodes including cAFM tip (Pt) and ITO. We approximate the molecular layer as an electron/hole tunneling barrier and attribute the resistance change caused by the magnetic field to the tunneling barrier modification. In the following, we employ a simplified one-dimensional barrier model and extract the barrier information by fitting the experimental I-V curves. Thereafter, we discuss the mechanism of barrier modification related to chirality and magnetic field.
The Simmon’s model [50] was modified to simulate a barrier tunneling problem as proposed in Ref. [13],
$I=g[(U+eV/2)e^{-A\sqrt{U+eV/2}}-(U-eV/2)e^{-A\sqrt{U-eV/2}}](1+\gamma V)$ (1)
where $U$ is the barrier height, $A$ is a parameter related to the effective mass and barrier width, $g$ is a parameter in the unit of conductance, and $\gamma$ characterizes the anisotropy between ±$\pm V$ (e.g., caused by the device asymmetry). We use Eq. (1) to fit our I-V data for a given magnetic field and extract $U,A,g,\gamma$. For an I-V curves corresponding to ±$\mathbf{\pm B}$ and B = 0 from the same Au NC, we use the same $A,g,\gamma$ but different $U$ in the fitting. We assume B only modifies the barrier $U$ but keeps other parameters unchanged. The fitted I-V curves were shown in Figures 2 and 4. We illustrate the barrier modulation and summarize the fitting barrier in Table 1 (see details in Extended Data Table. 4). One can find that independent fittings for L- and D-Au NCs give very similar barrier height ($U\approx 0.6$ eV) at zero field. The –B (+B) field enhances the barrier significantly to about 1.2 eV for L(D)-Au NCs. In contrast, –B (+B) reduces the barrier from about 1 eV to around 0.5 eV for Sm-L(D)-Au NCs.
To understand the opposite trend of barrier modification between L/D- and Sm-L/D-Au NCs, we performed further experiments to resolve their electronic structures. Considering the whole NC as a single particle, we measured the highest occupied molecular orbital (HOMO) and lowest unoccupied molecular orbital (LUMO) from the ionization energy and electron affinity by ultraviolet photoelectron spectroscopy (UPS) and low energy inverse photoemission spectroscopy (LEIPS), respectively, at B = 0 (see Figures 5c and 5d and Extended Data Figs. 17 and 18). For L/D-Au NCs, the LUMO energy (– 4.76 eV) is much closer to the Fermi energies of ITO and Pt compared to the HOMO energy. This indicates that electron rather than hole tunneling is dominant in the charge transport. Here, we can approximate the LUMO offset to electrode Fermi energy as the effective tunneling barrier, $0.1\sim 0.2$ eV as shown in Figure 5c. This barrier is smaller than the fitted value from the simplified model but still in the same order of magnitude. Because of different work functions between Pt (– 4.93 eV) and ITO (– 4.85 eV), the barrier profile (at zero bias) will be slightly tilted at zero bias as illustrated in Figure  5a. For Sm-L/D-Au NCs, however, the HOMO level (-5.01 eV) is closer to the electrode Fermi energies than LUMO, indicating the hole transport nature. The effective barrier for holes is also $0.1\sim 0.2$ eV here.
Known opposite tunneling carriers between L/D Au-NCs and Sm-L/D-Au NCs, we can rationalize their opposite barrier modulations by one scenario. Both the increase of electron barrier (L/D) and decrease of hole barrier (Sm-L/D) point to an up-shift of the NC electronic potential. In experiments, chiral Au NCs are all positively charged before and after sulfide modification, as revealed by the positive Zeta potential in Extended Data Fig. 19 and Extended Data Table 3. Thus, their molecule layer may prefer to trap extra electrons rather than holes. Then, electron-trapping increases both HOMO and LUMO energies of chiral NCs upon applying a selected magnetic field.
Our observation coincides with the charge-trapping theory proposed to understand CISS recently [13]. According to this theory, current flow tends to inject charge into the insulating molecular layer and the sign of charge (electron or hole) reverses if flipping magnetic field or chirality. The insulating layer may capture the charge in some charge-trapping centers which commonly exist in organic materials. As discussed above, the molecular shell can trap electrons. Therefore, induced electron-trapping increases the electronic potential in a selected magnetic field. In an opposite magnetic field, however, injected holes cannot be trapped in the shell layer and then hardly alter the electronic potential compared to the zero-field case. Consequently, the tunneling resistance is dramatically modified by a selected magnetic field but remains unchanged in the opposite field. Here, switching chirality exhibits a similar resistance response to the flipping of the magnetic field.
To summarize the possible mechanism, the chiral morphology of the Au NC core and the external magnetic field jointly drive electrons or holes into the capping molecule shell surrounding the Au NC. Then the insulating shell preferentially traps electrons rather than holes. When electrons are trapped in the shell layer, it increases the electronic potential of the nanocrystal. This change in potential modifies the tunneling barrier for charge transport through the nanocrystal. For electron tunneling (in L/D-Au NCs), the increased potential raises the barrier height and increases resistance. For hole tunneling (in Sm-L/D-Au NCs), the increased potential lowers the barrier height for holes and decreases resistance. The asymmetry arises because electron trapping only occurs for one magnetic field direction (depending on the chirality), while injected holes in the opposite field cannot be trapped. This mechanism further shows that switching chirality has a similar effect to flipping the magnetic field direction.
We note that this mechanism provides a qualitative understanding of our observations. A more quantitative description, incorporating microscopic material properties, will be the subject of future work. This model serves as a foundation for understanding the observed phenomena and guides further investigations. Nevertheless, the magnetochiral charge trapping model can provide helpful insights into chirality-driven phenomena including earlier CISS experiments. As predicted in Ref. [13], the chirality-driven charge trapping effect accounts magneto-transport with a magnetic electrode or only an external magnetic field, and can explain CISS MR and the unusual asymmetric MR in the same framework. As shown by the control experiments, chirality of Au NCs rather than chirality of surface molecules is decisive for MR. In chiral Au NCs, conducting electrons from the Au core feel B sensitively via Lorentz force and show cyclotron motion in the plane perpendicular to B (the orbital effect). When current flows, electrons move in a spiral pathway that can coincide or contradict the chiral potential from the NC morphology. We note that the magnetic length is about $l_{B}=25.6~{}\mathrm{nm}/\sqrt{B}=148$ nm for B = 0.03 T, which is smaller than NC sizes, supporting the cyclotron motion picture. As proposed in Ref. [13], chirality and magnetic field jointly induce charge trapping to the shell layer when current flows. For achiral Au NCs with chiral Cys shells, conducting electrons in Au feel the magnetic field but hardly sense the chiral potential of the insulating molecular shell, explaining the negligible change of resistance in this case. This mechanism also rationalizes why the previous CISS setup with chiral molecules sandwiched between two nonmagnetic electrodes shows unchanged resistance when applying a magnetic field. Because they are insulating, chiral molecules have no free electron to sense the magnetic field in such an experiment. Instead, they require a ferromagnetic electrode to contact chiral molecules so that chirality and magnetism coexist in the interface regime. Thus, we require that conducting free electrons feel both chiral geometry and magnetic field (or magnetism) simultaneously to manifest substantial changes in magneto-transport.
6 Perspectives and Conclusions
Our work opens up exciting future directions for enantiomer-selective modification of surface potentials on chiral NCs using external magnetic fields, with potential applications in controlling chemical and biological reactions. For instance, a negative magnetic field increases the electron potential energy for L-/Sm-L-Au NCs while leaving that of D-/Sm-D-Au NCs unchanged. This feature could enable fine-tuned regulation of surface reactions, such as redox processes and electrostatic interactions, via external magnetic fields. Such control could lead to the development of novel magnetically driven strategies for metal-catalyzed reactions [51] and specific enzymatic binding interactions. Furthermore, the magneto-enantiomer selection of chiral NCs demonstrates sensitivity to surface ligands, as evidenced by the opposite MR signs displayed by L-Au and Sm-L-Au NCs under the same magnetic field. This property suggests potential applications of these chiral NCs as a new class of indicators for distinguishing molecules with different binding affinities to chiral surfaces. Additionally, the unique responses of chiral NCs to magnetic fields could be harnessed for the controllable separation of molecular mixtures. By applying a designated magnetic field, it may be possible to selectively modify the surface properties of specific enantiomers, facilitating their separation from a mixture. These potential applications highlight the broad implications of our findings, spanning from fundamental surface chemistry to practical applications in catalysis, molecular recognition, and enantiomer separation. Future research in this direction could lead to innovative technologies leveraging the interplay between chirality, magnetism, and surface chemistry at the nanoscale.
In summary, we observed remarkable chirality-driven asymmetric MR in chiral Au NCs at the single nanoparticle level. This unusual MR exhibits distinct differences from CISS and EMCA, characterized by its asymmetric nature and absence of ferromagnetic electrodes. Notably, a simple sulfide modification of the chiral Au NCs enables switched positive and negative MR responses. We attribute the salient MR behavior to the asymmetric charge trapping influenced by both the magnetic field and chirality. Our work establishes the theory framework to understand different chirality-driven phenomena and paves a path to perform more quantitative calculations in the future. The implications of this unique MR mechanism extend beyond CISS and hold potential in the application of advanced crystalline electronic/spintronic devices and single-particle sensors. Our work indicates an exotic surface potential activation of chiral Au NCs and remote control of charge transfer in chiral Au NCs by magnetic fields. The charge transfer under a magnetic field can be selectively inhibited or enhanced by controlling their surface chemistry, providing a new strategy for manipulating enantiomer-specific chemical or biological reactions and the development of chemical sensors.
Methods
Synthesis of CPC-capped Au seeds. The CPC-capped Au seeds were synthesized according to procedures published elsewhere [52, 53], and they were used for the synthesis of different types of Au NCs. The major samples and their corresponding abbreviations in this report are listed in Extended Data Table 1.
Synthesis of chiral Au NCs. The synthesis of chiral Au NCs is realized through a Cys-directed growth method, which is based on the enantioselective interactions of L- or D-Cys with R- or S-$\{hkl\}$ chiral Au crystal facets ($h>k>l>0$[23, 30]. The R- or S-$\{hkl\}$ chiral Au crystal facet is defined as the clockwise or anticlockwise rotational progression among the (111), (100), and (110) microfacets around the kink atoms. In a typical synthesis, a growth solution was prepared by gently mixing 2.025 mL diluted water, 1.8 mL of 100 mM CPC and 100 mM CPB solution (CPC:CPB=1:3), 0.2 mL of 10 mM HAuCl4 solution, 1.4 mL of 0.1 M AA solution, 12.5 $\mu$L of 0.1 mM L- or D-Cys solution at 30 C. With the addition of Au seed solution, the growth of chiral NCs was started and continued at 30 C for 2 h. Then, the solution was centrifuged (6000 rpm, 3 min) and washed twice and then stored in 1 mM CTAB solution.
For the addition of 3, 6, 12, 24, 48, and 80 $\mu$L of seeds, the final sizes of the chiral Au NCs are 475.5±6.4 nm, 377.2±6.3 nm, 350.5±5.4 nm, 225.7±5.6 nm, 188.8±6.2, and 150.2±4.8 nm, respectively.
The CATB bilayer structure involves the adsorption of bromide anions on the Au surface, establishing an electrostatic interaction with the $\mathrm{(CH_{3})_{3}N^{+}}$ head group of the cetyltrimethylammonium cation ($\mathrm{CTA^{+}}$[33]. This arrangement leads to the outward orientation of the long hydrophobic alkane tail of $\mathrm{CTA^{+}}$, promoting the formation of a second layer through interactions between the alkane chains of other CTA+ cations. Consequently, the cationic $\mathrm{(CH_{3})_{3}N^{+}}$ groups remain on the external sides of the bilayer and the chiral Au NCs exhibit a positive charge.
Synthesis of achiral Au NCs from racemic Cys. Achiral Au NCs from racemic Cys were synthesized with the addition of 6.25 $\mu$L of 0.1 mM L-Cys and 6.25 $\mu$L of 0.1 mM D-Cys solution while keeping other parameters constant.
Synthesis of achiral Au NCs in the absence of Cys. A 0.2 mL of 10 mM HAuCl4 solution, 40 $\mu$L of 100 mM AA solution, and 80 $\mu$L of CPC-capped Au seeds solution were added to 4 mL of 50 mM CTAC solution at 30 C in sequence. The mixture was thoroughly mixed and undisturbed for 30 min. Then, the solution was centrifuged (6000 rpm, 3 min), washed twice and redispersed in water.
Modification of achiral Au NCs with L- or D-Cys. 80 $\mu$L of 100 mM L- or D-Cys solution was added into 2 mL of 4 mM the achiral Au NC solution (synthesized in the absence of Cys). The mixture was thoroughly mixed and incubated for 1 h at 30 °C. Then, the solution was centrifuged (6000 rpm, 3 min), washed once, and stored in 1 mM CTAB solution.
Modification of chiral Au NCs with sulfide. 1 mL of 2 mM $\text{Na}_{2}$ solution was added into 2 mL of the L- or D-Au NC solution (1 mM). The mixture was thoroughly mixed and incubated for 15 min at 30 °C. Then, the solution was centrifuged (6000 rpm, 3 min), washed once, and stored in 1 mM CTAB solution.
Characterization. Scanning electron microscopy (SEM) images were taken using a ZEISS Sigma 300 field-emission microscope equipped with an accelerating voltage of 20 kV. Extinction and circular dichroism (CD) spectra were obtained using a J-820 spectropolarimeter instrument (JASCO). Transmission electron microscope (TEM) were obtained from a Japan’s Hitachi H-600 TEM operating at 100 kV. Extinction spectra were collected by Shimadzu UV-1800 UV-vis spectrophotometer. X-ray photoelectron spectroscopy (XPS) spectra were performed with an ESCALAB MKII (VG Co., UK) spectrometer with an AlKa excitation source. Dynamic light scattering and zeta potential measurements were carried out on Malvern Zetasizer (Malvern Instrument Ltd, UK).
Preparation of individual Au NCs on the ITO. The indium tin oxide (ITO) substrates ($R=10\;\Omega$) were cleaned by sonication in ultrapure water ($18.2\;M\Omega\cdot\text{cm}$) and ethanol for 10 min, respectively. Thereafter, the ITO were washed with ultrapure water and dried with nitrogen. The Au NC solution (50 $\mu$L) was dropped on the ITO conductive surface. After natural drying, the Au NCs were dispersed on ITO substrates.
cAFM measurements. The images of Au NC on ITO substrates were measured using a NX 10 Complete AFM System (Park, Korea) with non-contact mode (NCM). The current-bias (I-V) curves of single Au NC were recorded in conducting atomic force microscopy (C-AFM) mode by applying the bias (from -1.0 V to 1.0 V) to the single Au NC at loading force of 10 nN. The Pt coated tip (HQ: XSC111, spring constant: $\sim$42 N/m) were used, and the I-V curves via the ITO/single Au NC/Pt junction were collected under external magnetic fields (300 Oe). For each type of Au NC, hundreds of I-V curves were collected from at least 20 individual Au NCs with more than three parallel samples. The log$\left|I\right|$ values at -1.0 V were extracted for statistics. All statistical histograms exhibited a normal distribution. The AFM images before and after I-V measurements display the same Au NCs morphologies, indicating the intact Au NCs during the I-V measurements.
For the measurements with varied loading forces, a force of 130 nN would cause damage to the sample, so the maximum force was limited to 100 nN. Various forces were applied in situ on the same Au NC, and 20-40 I-V curves were collected for each loading force. Under each magnetic field condition, 3-5 individual Au NCs were measured. While the absolute current values are slightly different among different Au NCs, the trends of current behavior and MR remain consistent.
Energy levels measurements. The ionization energy and electron affinity energy of Au NCs thin film on ITO substrate were obtained by ultraviolet photoelectron spectroscopy (UPS) and low energy inverse photoemission spectroscopy (LEIPS), respectively. UPS measurements were carried out using an ESCALAB 250Xi (Thermo Fisher, USA) spectrometer with He I line (21.22 eV) at a bias of -5.0 V. LEIPS measurements were performed on a customized ULVAC-PHI LEIPS instrument with Bremsstrahlung isochromatic mode.
Acknowledgements
The authors thank the financial support by the National Natural Science Foundation of China (Nos. 22374144 to W.N., 22374109 and 21974102 to G.C., 22072144 to F.W., 22102171 to Y.T.) and the Israel Science Foundation (ISF, No. 2932/21 to B.Y.).
Competing Interests
The authors declare no competing interests.